OpenStudy (anonymous):

how to simplify cos(2w)/cos(w)+sin(w)

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(2w)}{cos(w)+sin(w)}\quad ?\)

OpenStudy (anonymous):

yes how to simplify using trig identities

OpenStudy (jdoe0001):

http://www.mathwords.com/t/trig_identities.htm <--- look at the double-angle identities

OpenStudy (jdoe0001):

\(\bf \cfrac{cos(2w)}{cos(w)+sin(w)}\\ \quad \\ \quad \\ \textit{notice that }\quad \color{blue}{cos(2\theta)=cos^2(\theta)-sin^2(\theta)}\qquad thus\\ \quad \\ \cfrac{cos(2w)}{cos(w)+sin(w)}\implies \cfrac{cos^2(w)-sin^2(w)}{cos(w)+sin(w)}\\ \quad \\ \textit{now recall that }\quad \color{blue}{a^2-b^2 = (a-b)(a+b)}\qquad thus\\ \quad \\ \cfrac{cos^2(w)-sin^2(w)}{cos(w)+sin(w)}\implies \cfrac{[cos(w)-sin(w)][cos(w)+sin(w)]}{cos(w)+sin(w)}\)

OpenStudy (jdoe0001):

and you can see what cancels out there

OpenStudy (anonymous):

thank you thank you thank you!!!!

OpenStudy (jdoe0001):

yw